A random variable, x, has a normal distribution with μ = 15.9 and σ = 2.80. Determine a value, x0, so that:

a. P(x>x0) = 0.05

x0 = ______ . (Round to one decimal place as needed.)

b. P(x≤x0) = 0.975

x0 = ______ . (Round to one decimal place as needed.)
Answer #1

Solution :

Using standard normal table,

a.

P(Z > z) = 0.05

1 - P(Z < z) = 0.05

P(Z < z) = 1 - 0.05

P(Z < 1.65) = 0.95

z = 1.65

Using z-score formula,

x = z * \sigma + \mu

x = 1.65 * 2.80 + 15.9 = 20.5

b.

P(Z \leq z) = 0.975

P(Z \leq 1.96) = 0.975

z = 1.96

Using z-score formula,

x = z * \sigma + \mu

x = 1.96 * 2.80 + 15.9 = 21.4